OpenStudy (anonymous):

Consider the following pair of equations: y = x + 4 y = −2x − 2 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

HanAkoSolo (jamierox4ev3r):

by doing substitution, you are effectively combining the two equations together. Therefore, we need to change the easier equation so that it would be equal to x

HanAkoSolo (jamierox4ev3r):

The key thing here is that if we set an equation equal to x, then we can plug in that equation in place of the x in the second equation.

HanAkoSolo (jamierox4ev3r):

Following me so far? @Taya1431

HanAkoSolo (jamierox4ev3r):

So right now, I want you to set y=x+4 so that it equals x So solve for x.

OpenStudy (anonymous):

x=y-4

HanAkoSolo (jamierox4ev3r):

yes. So now that we know x is equal to y-4, we can insert this into the other equation. We are "substituting" the variable for an actual definition. y=-2(y-4)-2 ^ Now we only have one variable, so we can solve for y

HanAkoSolo (jamierox4ev3r):

Do you see what I did there?

OpenStudy (anonymous):

so you subtract the 2?

HanAkoSolo (jamierox4ev3r):

you replace the "x" in the second equation with "y-4" and before you subtract the two, distribute the -2 into the parentheses. Can you do that for me?

OpenStudy (anonymous):

uh.. like this? ~ 2=2-4

HanAkoSolo (jamierox4ev3r):

No, you still have the two...Here, let me show you what I mean by distribution. And sorry for the late replies, my internet is being really weird

OpenStudy (anonymous):

its cool x3

HanAkoSolo (jamierox4ev3r):



OpenStudy (anonymous):

oh..

HanAkoSolo (jamierox4ev3r):

So we went from this: y=-2(y-4)